OpenStudy (anonymous):

A 1.5 kg rolling pin rolls along a countertop at 6.0 m/s. It collides with and sticks to a 200 g lump of dough. The doughy rolling pin continues along the countertop. What is the final speed of the doughy rolling pin?

OpenStudy (shamim):

Use m1*u1+m2*u2=m1*v1+m2*v2 m1*u1+m2*u2=(m1+m2)v v=?

OpenStudy (shamim):

Response plz!!"